# Hidden Population III Descendants in Ultra-Faint Dwarf Galaxies

### v1

@@ 4.3 Does Boötes I host stars mono-enriched by Pop III SNe?
Here, we define ``````` `mono-enriched”  stars as those that have originated from gas enriched by exactly one Pop III supernova. Fig. 5 shows that in Boötes I we can potentially find mono-enriched stars which cover wide ranges in [Fe/H] and [C/Fe], and are enriched by different types of Pop III SNe. In line to what we discussed already, for $\rm[C/Fe]>+0.7$ and $\rm[Fe/H]\lesssim-3.5$ we find predominantly mono-enriched stars imprinted by faint, cc, high energy. Mono-enriched CEMP stars are rare $\sim 5\%$ with respect to the total of CEMP stars in Boötes I, and $\sim 5\%$ with respect to CEMP stars predominantly enriched by Pop III. On the other hand, mono-enriched stars by hypernovae are typically found at $\rm[C/Fe]<+0.7$ and $\rm[Fe/H]\gtrsim-3.5$. Their fraction is maximum for $\rm[C/Fe]<-0.5$ representing $10\%$ with respect the total number of stars in this range.
@@ 5 Progenitors of individual Boötes I stars
We can now use the predictions of our chemical evolution model to uncover hidden Pop III descendants in Boötes I. To this end, we analyze the measure abundance patterns of Boötes I stars available in the literature, for which at least five chemical abundance ratios have been measured. To identify the progenitors of Boötes I stars, we compared the abundance patterns of simulated stars that survive until today with those of observed stars. Firstly we pre-selected models that have [Mg/Fe] and [Fe/H] abundance ratios that are consistent with the observed values, i.e., within the error bars. Then, among all models reproducing [Mg/Fe] and [Fe/H], we found those that best fit all the measured chemical elements by minimizing the $\chi^{2}$ between the simulated and observed [X/Fe] ratios. Finally, after selecting the best fits, we retrieve from the simulation the propertties of the SNe responsible for enriching the birth cloud from which these stars originated. Note that the [C/Fe] ratio of the observed stars in Boötes I have been corrected for the effect of evolutionary status by exploiting the online tool [^4] http://vplacco.pythonanywhere.com/ presented in Placco et al. (2014), while the abundance ratios are not corrected for 3D and Non-Local Thermodynamic Equilibrium (NLTE) effects.
We categorized Boötes I stars into the following groups, according to their progenitors: i) Hidden Pop III descendants, long-lived stars enriched by Pop III SNe. ii) Ambiguous stars, whose abundance patterns are consistent with a predominant enrichment from both Pop III and/or Pop II stars, making it challenging to distinguish between these two scenarios; and iii) Pop II descendants, whiich are primarily enriched by the normal Pop II stellar population. Below, we will concentrate on the truly/ambiguous Pop III descendants, i.e., groups i) and ii), while in Appendix B we discuss group iii).
@@ 5.1 Hidden Pop III descendants
Three stars in Boötes I have abundance patterns consistent with a $100\%$ enrichment from Pop III SNe. When we refer to a $100\%$ enrichment by Pop III SNe, we are indicating that the star formed in an environment enriched solely by Pop III SNe. However, this does not imply that the star has been enriched by a single supernova, i.e. that the star is a mono-enriched Pop III descendant. For this reason, in the following, we will differentiate between multi- and mono-enriched Pop III descendants. The comparison between the measured abundance patterns and the best-fits ones are shown in Fig.6. Additionally, we show the abundance pattern of a star enriched by multiple Pop II stellar populations (100% Pop II enriched star) with the chi-square closest to the minimum.
@@ 5.1.1 Mono-enriched Pop III descendants
In Boötes I we uncover one mono-enriched Pop III descendant, i.e. a star imprinted by a single Pop III SNe:
Boo-130: The best fit shows that Boo-130 is mono-enriched by a hypernova with $m_{\star}=50\>\rm{M_{\odot}}$ and $\rm E_{SN}=10\times 10^{51}$ erg (Fig.6). Note that in this case, [Zn/Fe] is the key abundance ratio to discriminate between an enrichment driven by Pop III and Pop II stellar populations.
Figure 6: Comparison between the measured abundance patterns of three Boötes I stars classified as hidden Pop III descendants (red star symbols) and our best-fits (squares). From top to bottom we show the data of CEMP Boo-119 star (Gilmore et al., 2013), two C-normal stars Boo-94 (Ishigaki et al., 2014) and Boo-94 (Gilmore et al., 2013) and C-low star Boo-130 (Ishigaki et al., 2014). In grey the predicted abundance patter for an enrichment driven by Pop II stellar population with the chi-square closest to the minimum.
@@ 5.1.2 Multi-enriched Pop III descendants
Among the identified three Pop III descendants, we found that two of them are multi-enriched by Pop III SNe with different masses and/or different energies:
Boo-119: The measured abundance pattern of the CEMP ($\rm[C/Fe]\approx+2.5$) star Boo-119 (Gilmore et al., 2013) is consistent with $100\%$ faint Pop III SNe enrichment (top panel of Fig.6). However, according to our predictions, Boo-119 was enriched by two faint Pop III SNe with similar progenitor masses: i) $\rm E_{SN}=0.3\times 10^{51}\>erg$ and $m_{\star}=55\>\rm{M_{\odot}}$ (50%); and ii) $\rm E_{SN}=0.6\times 10^{51}\>erg$ and $m_{\star}=58\>\rm{M_{\odot}}$ (50%). As we can see, our predicted abundance pattern nicely reproduces that observed except for [Ti/Fe], which is commonly underestimated in models compared to observations.
Conversely, the best-fit among all the abundance patterns solely driven by a Pop II SN enrichment, fails to reproduce the observed [C/Fe] value, which is lower by $>1$ dex, as well as the [Ca/Fe] and [Cr/Fe] ratios, which exhibit a discrepancy of $\sim 0.5$ dex.
Boo-94: The measured abundance pattern of the Boo-94 star is consistent with $100\%$ enrichment by three Pop III SNe with different energies ($\sim 24\%$ by high-energy, $\sim 76\%$ by cc) and masses: two ccSNe with $\rm E_{SN}=1.2\times 10^{51}\>erg$, and $\rm E_{SN}=1.5\times 10^{51}\>erg$, and masses respectively equal to $m_{\star}=26.5\>\rm{M_{\odot}}$, and $m_{\star}=32\>\rm{M_{\odot}}$; and one high-energy SNe with $\rm E_{SN}=3.0\times 10^{51}\>erg$, $m_{\star}=27\>\rm{M_{\odot}}$. Also for this star, the predicted abundance pattern of a $100\%$ Pop II enrichment fails to accurately reproduce the observed abundances in [C/Fe], [Na/Fe], [Ca/Fe], [Ti/Fe] and [Zn/Fe], see Fig.6.
@@ 5.2 Ambiguous stars
Within the analyzed Boötes I stars, we have identified a subset that we now classify as ``````` `ambiguous”. An example of this is displayed in Fig.7, where we show the comparison between the measured abundance pattern of Boo-1137 (Norris et al. 2010) and the simulated one that produce the best fit. For this star, the minimum $\chi^{2}\approx 4.0$ is obtained with an enrichment mainly driven by Pop III SNe. However, when plotting the predicted abundance pattern for a 100% Pop II enrichment (with a > 80% contribution by AGB stars) with a chi-square value very close to the minimum ($\chi^{2}\approx 4.5$), we observe that these two abundance patterns appear strikingly similar. Specifically, the two predicted abundance patterns are almost indistinguishable in the Na, Mg, Al, Ca, Sc, Ti, Cr, Mn, Co, and Ni abundances, while they exhibit differences of $>1\,\rm dex$ in C, N, O, and Zn. It is worth noting that this degeneracy could be resolved with the measurement of [N/Fe]. However, nitrogen is both challenging to measure accurately and the surface abundance strongly depends on the stellar evolutionary stage, due to internal stellar mixing processes (e.g. Gratton et al., 2000; Placco et al., 2014).
Why an enrichment driven by 100% Pop III is so similar to 100% Pop II SNe? The reason lies in the number of Pop III SNe progenitors contrtibuting to the ISM pollution. Indeed, upon retrieving the progenitors, we found that this abundance pattern is the result of enrichment from multiple Pop III SNe (>10), with different masses and energies. Consequently, the distinctive features of individual Pop III SNe are lost resulting in an abundance pattern very similar to that driven by Pop II stars (see Fig.7).
Figure 7: Comparison between the measured abundance pattern of the star Boo-1137 classified as ambigous (grey star symbols, Norris et al. 2010) and our best-fit models (squares) for an enrichment driven by Pop II (green) and by Pop III stellar (red) populations. Note that the Pop II abundance pattern is mainly driven by the contribution of Pop II AGB stars.
@@ 6 Probing Pop III Supernova properties
According to our model, the three Pop III SNe descendants have been enriched by a total of two faint, two cc, one high-energy SN and one hypernovae, i.e. four out of six Pop III progenitors are low-energy SNe. Despite the limited statistics, our findings seem to suggest that Boötes I preferentially retains the chemical elements released by low-energy Pop III SNe. To further explore this aspect, in Fig. 9 we evaluate the percentage of metals, coming from different SNe types, that are locked in stars surviving until $z=0$. Given that Boötes I is currently a gas-free galaxy, this percentage also reflects the metals retained by its gravitational potential well throughout its evolution.
@@ 7 Discussion
Among the 42 Boötes I stars with measured chemical abundance ratios, three are consistent with being direct Pop III descendants. Notably, two of these three stars are multi-enriched by Pop III SNe with different energies and masses (see Sec.5.1). On the other hand, our results show that if the gas has been enriched by more than 10 Pop III SNe, the resulting abundance pattern does not display the distinctive chemical features of Pop III descendants. Instead, it becomes indistinguishable from that produced by Pop II stars (see Sec.5.2). Boo-1137 is a perfect example of this case (see Fig.7), where we cannot discriminate between enrichment driven by a Pop III or Pop II stellar population. This star was also analyzed by Welsh et al. (2023), who found that Boo-1137 had been enriched by a single Pop III progenitor with $m_{\star}=(18-27)\>\rm{M_{\odot}}$. It is important to note that to achieve this result, the authors used only five chemical elements (C, O, Al, Si, Fe), while for Boo-1137, there are 11 other measured chemical elements in the literature that we used to determine the best-fit model. We can speculate that the different interpretation of the origin of this Boo-1137 is due to the number of chemical abundance ratios used in the analysis. This discrepancy emphasizes the importance of employing a wide range of chemical elements (> 5) to accurately interpret data and to give constraints on the initial metallicity, mass and explosion energy of the progenitor stars imprinting the birth environment.
@@
For direct Pop III descendants we infer the mass of their progenitors that is in mass range $m_{\star}=[20-60]\>\rm{M_{\odot}}$. These results are partially in agreement with the results of Ishigaki et al. (2018). The authors found that the abundance patterns of extremely metal-poor stars ($\rm[Fe/H]<-3$) are predominantly best-fitted by SNe yields with $m_{\star}<40\>\rm{M_{\odot}}$ and that more than than half of the stars are best-fitted by the $m_{\star}=25\>\rm{M_{\odot}}$ hypernova $(E=10\times 10^{51}\rm erg)$ models. Finally they conclude that the masses of Pop III stars responsible for the first metal enrichment are predominantly $m_{\star}<40\>\rm{M_{\odot}}$. On the other hand, our results suggest that the masses of Pop III stars responsible of the enrichment are higher, up to $\approx 60\>\rm{M_{\odot}}$. Note that Ishigaki et al. (2018) use different sets of stellar yields (see their Sec.2) and they have, for $m_{\star}>25\>\rm{M_{\odot}}$, only two stellar model yields for stars with $m_{\star}=40\>\rm{M_{\odot}}$ and $m_{\star}=100\>\rm{M_{\odot}}$ (see Table 1 in Ishigaki et al., 2018). While we used the stellar yields of Heger & Woosley (2010) that have more stellar models within this range.

### v2

@@ 5.3 Does Boötes I host stars mono-enriched by Pop III SNe?
Here, we define ``````` `mono-enriched”  stars as those that have originated from gas enriched by exactly one Pop III supernova. Fig. 5 shows that in Boötes I we can potentially find mono-enriched stars which cover wide ranges in [Fe/H] and [C/Fe], and are enriched by different types of Pop III SNe. In line to what we discussed already, for $\rm[C/Fe]>+0.7$ and $\rm[Fe/H]\lesssim-3.5$ we find predominantly mono-enriched stars imprinted by faint, cc, and high energy SNe. Mono-enriched CEMP stars are rare $\sim 5\%$ with respect to the total of CEMP stars in Boötes I, and $\sim 5\%$ with respect to CEMP stars predominantly enriched by Pop III. On the other hand, mono-enriched stars by hypernovae are typically found at $\rm[C/Fe]<+0.7$ and $\rm[Fe/H]\gtrsim-3.5$. Their fraction is maximum for $\rm[C/Fe]<-0.5$ representing $10\%$ with respect to the total number of stars in this range.
@@ 6 Progenitors of individual Boötes I stars
We can now use the predictions of our chemical evolution model to uncover hidden Pop III descendants in Boötes I. To this end, we analyzed the measured abundance patterns of Boötes I stars available in the literature, for which at least five chemical abundance ratios [X/Fe] have been measured.
To identify the progenitors of Boötes I stars, we compared the abundance patterns of simulated stars that survive until today with those of observed stars. Firstly we pre-selected models that have [Fe/H] consistent with the observed values, i.e., within the error bars. Then, among all models reproducing [Fe/H], we found those that best fit all the measured chemical elements by minimizing the reduced chi-squared, $\chi^{2}$, between the simulated and observed [X/Fe] ratios:
$\chi^{2}=\sum_{i=1}^{N_{\text{obs}}}\frac{\left([\text{X/Fe}]_{\text{obs},i}-[\text{X/Fe}]_{\text{sim},i}\right)^{2}}{\sigma_{[\text{X/Fe}]\text{obs},i}^{2}+\sigma_{[\text{X/Fe}]\text{sim},i}^{2}}$ (5)
where $\rm[X/Fe]_{\text{obs},i}$ represents the observed abundance with its corresponding uncertainty $\sigma_{\rm[X/Fe]_{\text{obs},i}}$, while$\rm[X/Fe]_{\text{sim},i}$ denotes the simulated abundance for the same element, with an assumed uncertainty of $\sigma_{\rm[X/Fe]_{\text{sim},i}}=0.2$. Finally, after selecting the best fits, we retrieve from the simulation the properties of the SNe responsible for enriching the birth cloud from which these stars originated. Note that the [C/Fe] ratio of the observed stars in Boötes I have been corrected for the effect of evolutionary status by exploiting the online tool [^5] http://vplacco.pythonanywhere.com/ presented in Placco et al. (2014), while the abundance ratios are not corrected for 3D and Non-Local Thermodynamic Equilibrium (NLTE) effects.
We categorized Boötes I stars into the following groups, according to their possible progenitors: i) Hidden Pop III descendants, long-lived stars consistent with $100\%$ enrichment by Pop III SNe. ii) Ambiguous stars, whose abundance patterns are consistent with a predominant enrichment from both Pop III and/or Pop II stars, making it challenging to distinguish between these two scenarios; and iii) Pop II descendants, which are primarily enriched by the normal Pop II stellar population. Below, we will concentrate on the truly/ambiguous Pop III descendants, i.e., groups i) and ii), while in Appendix B we discuss group iii).
@@ 6.1 Hidden Pop III descendants
Three stars in Boötes I have abundance patterns consistent with a $100\%$ enrichment from Pop III SNe, therefore making them good candidates for being Pop III descendants.. When we refer to a $100\%$ enrichment by Pop III SNe, we are indicating that the star formed in an environment enriched solely by Pop III SNe. However, this does not imply that the star has been enriched by a single supernova, i.e. that the star is a mono-enriched Pop III descendant. For this reason, in the following, we will differentiate between multi- and mono-enriched Pop III descendants. The comparison between the measured abundance patterns and the best-fits ones are shown in Fig.6. The figure also displays the two models with $\chi^{2}$ closest to the minimum, which in all cases also have 100% Pop III enrichment. Additionally the best fit for 100% Pop II enrichment is shown, but this has significantly higher $\chi^{2}$ in all three cases.
@@ 6.1.1 Mono-enriched Pop III descendants
In Boötes I we uncover one potential mono-enriched Pop III descendant , i.e. a star imprinted by a single Pop III SNe:
Boo-130: The best fit indicates that Boo-130 is mono-enriched by a Pop III hypernova with $m_{\star}=64\>\rm{M_{\odot}}$ and $\rm E_{SN}=5\times 10^{51}$ erg (Fig.6). Note that in this case, [Zn/Fe] is the key abundance ratio to discriminate between an enrichment driven by Pop III and Pop II stellar populations. Importantly, the models with $\chi^{2}$ values closest to the minimum also point to a $100\%$ Pop III enrichment, in this case by multi-enrichment. This strengthens the hypothesis that Boo-130 is a strong candidate for being a Pop III descendant. While some degeneracy remains with other models featuring comparable $\chi^{2}$-values, Boo-130 can be considered a descendant of Pop III stars if we rely on the current model constraints.
Figure 6: Measured abundance patterns of three Boötes I stars classified as hidden Pop III descendants (red star symbols) compared to our best fits (colored lines). From top to bottom we show the data of the CEMP-no Boo-119 star (Lai et al., 2011; Gilmore et al., 2013); the CEMP-no star Boo-94 (Ishigaki et al., 2014; Gilmore et al., 2013); and the C-normal star Boo-130 (Ishigaki et al., 2014). The second (filled circles, dashed lines) and third (open circles, dotted lines) best fits are also shown in gray along with the Pop II enrichment scenario with the lowest $\chi^{2}$ (squares, solid lines).
@@ 6.1.2 Multi-enriched Pop III descendants
Among the three identified possible Pop III descendants, we found that two of them are multi-enriched by Pop III SNe with different masses and/or energies:
Boo-119: The measured abundance pattern of the CEMP ($\rm[C/Fe]\approx+2.5$) star Boo-119 (Lai et al. 2011 for [C/Fe], Gilmore et al. 2013 for rest) is consistent with $100\%$ faint Pop III SNe enrichment (top panel of Fig.6). However, according to our predictions, Boo-119 was enriched by two faint Pop III SNe with similar progenitor masses: i) $\rm E_{SN}=0.3\times 10^{51}\>erg$ and $m_{\star}=43\>\rm{M_{\odot}}$ (50%); and ii) $\rm E_{SN}=0.3\times 10^{51}\>erg$ and $m_{\star}=55\>\rm{M_{\odot}}$ (50%). As we can see, our predicted abundance pattern nicely reproduces that observed except for [Ti/Fe], which is commonly underestimated in models compared to observations.
Conversely, the best fit among all the abundance patterns solely driven by a Pop II SN enrichment, fails to reproduce the observed [C/Fe] value, which is lower by $>1$ dex, as well as the [Mg/Fe] and [Na/Fe] ratios, which exhibit a discrepancy of $\gtrsim 0.8$ dex. Note that the other models with $\chi^{2}$-values closest to the minimum also predict an enrichment driven by $100\%$ Pop III SNe, albeit with different contributions from different Pop III SNe types (Fig.6).
Boo-94: The measured abundance pattern of the Boo-94 star is consistent with $100\%$ enrichment by three Pop III SNe with different energies ($\sim 39\%$ by faint, $\sim 36\%$ by cc, $\sim 25\%$ by high-energy) and masses: with $\rm E_{SN}=[0.3,1.2,3]\times 10^{51}\>erg$ and masses respectively equal to $m_{\star}=[24,30,35]\>\rm{M_{\odot}}$. Also for this star, the predicted abundance pattern of a $100\%$ Pop II enrichment fails to accurately reproduce the observed abundances , in particular in regards to in [C/Fe], [Na/Fe] and [Zn/Fe], see Fig.6.
@@ 6.2 Ambiguous stars
Within the analyzed Boötes I stars, we have identified a subset that we now classify as ``````` `ambiguous”. An example of this is displayed in Fig.7, where we show the comparison between the measured abundance pattern of Boo-1137 (Norris et al. 2010) and the simulated one that produce the best fit. For this star, the minimum $\chi^{2}\approx 9$ is obtained with an enrichment mainly driven by Pop III SNe. However, when plotting the predicted abundance pattern for a 100% Pop II enrichment with a $\chi^{2}$-value very close to the minimum ($\chi^{2}\approx 12$), we observe that these two abundance patterns appear strikingly similar. Specifically, the two predicted abundance patterns are almost indistinguishable in the C, Al, Si, Ca, Sc, Ti, Cr, Mn, Co, Ni, Zn abundances, while they exhibit differences of $>1\,\rm dex$ in N, O, and Na , for which no measurements are available. It is worth noting that this degeneracy could be resolved with the measurement of [N/Fe]. However, nitrogen is both challenging to measure accurately and the surface abundance strongly depends on the stellar evolutionary stage, due to internal stellar mixing processes (e.g. Gratton et al., 2000; Placco et al., 2014).
Why an enrichment driven by 100% Pop III is so similar to 100% Pop II SNe? The reason lies in the number of Pop III SNe progenitors contributing to the ISM pollution. Indeed, upon retrieving the progenitors, we found that this abundance pattern is the result of enrichment from multiple Pop III SNe (>10), with different masses and energies. Consequently, the distinctive features of individual Pop III SNe are lost resulting in an abundance pattern very similar to that driven by Pop II stars (see Fig.7).
Figure 7: Comparison between the measured abundance pattern of the star Boo-1137 classified as ambigous (grey star symbols, Norris et al. 2010) and our best-fit models (squares) for an enrichment driven by Pop II (green) and by Pop III stellar (red) populations.
@@ 7 Probing Pop III Supernova properties
Our analysis strongly suggests that these stars are compelling candidates for being genuine Pop III descendants. However, as illustrated in Fig.6, a clear degeneracy remains: different combinations of Pop III SNe types, with varying progenitor masses and explosion energies, can produce equally good fits to the observed abundance patterns. While chi-square minimization identifies the best-fitting models, it is important to note that multiple solutions exist within a narrow range of $\chi^{2}$ values. Here the aim is to introduce a method to infer the physical properties of the Pop III progenitors, such as their masses and explosion energies, which shaped the chemical abundances observed in these stars. Naturally, obtaining more robust constraints will require additional measurements of chemical elements, which could help break the degeneracies and better differentiate between competing scenarios. Nonetheless, with the currently available data, we provide the most reliable estimates possible.
7.1 Metal Retention in Boötes I
According to our model, the three Pop III SNe descendants have been enriched by a total of three faint, one cc, one high-energy SN and one hypernovae, i.e. four out of six Pop III progenitors are low-energy SNe. Despite the limited statistics, our findings seem to suggest that Boötes I preferentially retains the chemical elements released by low-energy Pop III SNe. To further explore this aspect, in Fig. 9 we evaluate the percentage of metals, coming from different SNe types, that are locked in stars surviving until $z=0$. Given that Boötes I is currently a gas-free galaxy, this percentage also reflects the metals retained by its gravitational potential well throughout its evolution.
@@ 8 Discussion
Among the 54 Boötes I stars with measured chemical abundance ratios, three are consistent with being direct Pop III descendants. Notably, two of these three stars are multi-enriched by Pop III SNe with different energies and masses (see Sec.6.1). On the other hand, our results show that if the gas has been enriched by more than 10 Pop III SNe, the resulting abundance pattern does not display the distinctive chemical features of Pop III descendants. Instead, it becomes indistinguishable from that produced by Pop II stars (see Sec.6.2). Boo-1137 is a perfect example of this case (see Fig.7), where we cannot discriminate between enrichment driven by a Pop III or Pop II/I stellar population. This star was also analyzed by Welsh et al. (2023), who found that Boo-1137 had been enriched by a single Pop III progenitor with $m_{\star}=(18-27)\>\rm{M_{\odot}}$. It is important to note that to achieve this result, the authors used only five chemical elements (C, O, Al, Si, Fe), while for Boo-1137, there are 11 other measured chemical elements in the literature that we used to determine the best-fit model. We can speculate that the different interpretation of the origin of this Boo-1137 is due to the number of chemical abundance ratios used in the analysis. This discrepancy emphasizes the importance of employing a wide range of chemical elements (> 5) to accurately interpret data and to give constraints on the initial metallicity, mass and explosion energy of the progenitor stars imprinting the birth environment.
It is worth noting that all our results are based on our fiducial model, which adopts a flat EDF for PopIII stars and the PopII stellar yields from Limongi & Chieffi (2018). To test the robustness of our interpretation, we investigated whether the abundance patterns of our PopIII candidates could also be reproduced using alternative PopII enrichment scenarios. Specifically, we employed the yield tables from Nomoto et al. (2013) at a fixed metallicity of $\log(Z/Z_{\odot})=-3$, considering both standard core-collapse SNe and hypernovae. For each of our three Pop III-only candidates (Boo-119, Boo-130, and Boo-94), we computed the expected abundance patterns assuming enrichment from either a single Pop II SN or a combination of two Pop II progenitors with different masses. As shown in Fig.12 of Appendix C, adopting these alternative PopII sets of yields, we obtain best-fit models with higher chi-squared values compared to the PopIII-only enrichment scenarios, i.e., our fiducial model. This supports the interpretation that these stars are robust candidates for being genuine Pop III descendants.
@@
For direct Pop III descendants we infer the mass of their progenitors that is in mass range $m_{\star}=[20-65]\>\rm{M_{\odot}}$. These results are partially in agreement with the results of Ishigaki et al. (2018). The authors found that the abundance patterns of extremely metal-poor stars ($\rm[Fe/H]<-3$) are predominantly best-fitted by SNe yields with $m_{\star}<40\>\rm{M_{\odot}}$ and that more than half of the stars are best-fitted by the $m_{\star}=25\>\rm{M_{\odot}}$ hypernova $(E=10\times 10^{51}\rm erg)$ models. Finally, they conclude that the masses of Pop III stars responsible for the first metal enrichment are predominantly $m_{\star}<40\>\rm{M_{\odot}}$. On the other hand, our results suggest that the masses of Pop III stars responsible of the enrichment are higher, up to $\approx 65\>\rm{M_{\odot}}$. Note that Ishigaki et al. (2018) use different sets of stellar yields (see their Sec.2) and they have, for $m_{\star}>25\>\rm{M_{\odot}}$, only two stellar model yields for stars with $m_{\star}=40\>\rm{M_{\odot}}$ and $m_{\star}=100\>\rm{M_{\odot}}$ (see Table 1 in Ishigaki et al., 2018). While we used the stellar yields of Heger & Woosley (2010) that have more stellar models within this range.